OpenStudy (anonymous):

Capacitors C1 = 6.44 uF and C2 = 2.45 uF are charged as a parallel combination across a 241 V battery. The capacitors are disconnected from the battery and from each other. They are then connected positive plate to negative plate and negative plate to positive plate. Calculate the resulting charge on capacitor C1. Calculate the resulting charge on capacitor C2. I thought that electricity was making sense to me, I am evidently mistaken.

OpenStudy (anonymous):

Compute charge Q = C V for C1 and C2. Note that Q1/Q2 = C1/C2 = 92/35. In step 2 plus and minus charges cancel each other partially. The resulting charge Q1-Q2 is distributed again 92:35 across the capacitors. hth!

OpenStudy (anonymous):

OK, I did Q=CV, I got 6.44*10^-8C for Q1 and 1.02*10^-8C for Q2. I don't know what to do now to get my answers. I tried these answers but I was wrong.

OpenStudy (anonymous):

You should check your multiplications! Q1 = C1 V = 6.44*10^-6 * 241 = 1.552*10^-3 C. Q2 = C2 V. Then deltaQ=Q1-Q2....